A particle is fired at 22 degrees with respect to an incline and then hits a point on the incline as label. If the initial velocity is 15 m/s, a. find the time it takes for the projectile to hit the incline plane. b. find the distance from launching point to the landing point.


1
Expert's answer
2021-05-30T15:46:25-0400

a)

"t=\frac{2v\sin\alpha}{g}=1.15~s,"

b)

"l=\frac{v^2\sin2\alpha}{g}=16~m."